What is the maximum amount of weight that could be pulled up an inclined plane



if the force is limited to 50.75 pounds, the inclined plane is 12.25 feet long, and the



height is 54 inches?

1
Expert's answer
2021-11-14T17:17:10-0500

In order to pull the weight up an inclined plane, the following condition must satisfied:


"F_{pull}>Wsin\theta=W\dfrac{h}{L}."

Let's find the maximum amount of weight:


"W_{max}=\dfrac{F_{pull}L}{h},""W_{max}=\dfrac{50.75\ lb\times12.25\ ft}{54\ in\times\dfrac{1\ ft}{12\ in}}=138\ lb."